OpenStudy (anonymous):

For which operations is the set -1,0,1 closed? Choose all answers that are correct A: addition B: division C: multiplication D: subtraction

OpenStudy (anonymous):

You can select any number from that set and by adding, them if you get a number which is there in set too, you will say for Addition, the set is closed.. :)

OpenStudy (anonymous):

\(1 + 1 = ??\)

OpenStudy (mayankdevnani):

@Passion567 first of all,do you know what is the meaning of closed interval ?

OpenStudy (anonymous):

no

OpenStudy (mayankdevnani):

it means that when you apply any mathematical property on numbers,if the answer would come only and only -1,0,1 ,then we can say that the result(answer) is closed in the interval

OpenStudy (mayankdevnani):

understood? @Passion567

OpenStudy (anonymous):

somewhat

OpenStudy (mayankdevnani):

so for option A) if you add,then you will get :- -1+0=-1 1+0=1 -1+1=0 Can we say that option A is closed in the set?

OpenStudy (mayankdevnani):

could you get that ? @Passion567

OpenStudy (anonymous):

no i have no clue what this is i do homeschool and they dont teach us about it they just give us test and make us take them

OpenStudy (anonymous):

@mayankdevnani you need to check it once again, If I will do \(1+1\), then result will not be in the given interval, sorry, but I am here in doubt.. :(

OpenStudy (mayankdevnani):

well, @Passion567 see the result from option A is closed because the results are as follows:--1,0,1 and this result is include in the set which is given in the question. So,option A is correct

OpenStudy (mayankdevnani):

likewise,check all the other options

OpenStudy (mayankdevnani):

now @waterineyes you tried option D) subtraction See :- -1-(0)=-1---------->CLOSED 1-(0)=1------------->CLOSED -1-(1)=-2------------>OPEN(NOT CLOSED)

OpenStudy (mayankdevnani):

hope you understand. @Passion567 @waterineyes

OpenStudy (anonymous):

ok thank you

OpenStudy (mayankdevnani):

so option D is not our answer

OpenStudy (mayankdevnani):

so option D is not our answer :)

OpenStudy (mayankdevnani):

WELCOME :) @Passion567

OpenStudy (anonymous):

In my knowledge, option A ie addition is not correct one.. :)

OpenStudy (mayankdevnani):

why ???

OpenStudy (anonymous):

ok why

OpenStudy (mayankdevnani):

option A satisfies all the conditions

OpenStudy (anonymous):

I mean can't we add 1 with 1 itself?

OpenStudy (mayankdevnani):

why ???? only +1 and -1 is given in the question ?

OpenStudy (mayankdevnani):

You can try one number only once

OpenStudy (anonymous):

http://mathforum.org/library/drmath/view/52422.html

OpenStudy (anonymous):

1 and 1 both are in the set.. :)

OpenStudy (mayankdevnani):

yeah...you are right !!!

OpenStudy (anonymous):

but I am still in doubt.. :P

OpenStudy (mayankdevnani):

@Passion567 @waterineyes is right We have to include all the numbers

OpenStudy (mayankdevnani):

no,i got it @waterineyes

OpenStudy (anonymous):

Wait, I will give one other link which will contradict me and prove you are right, and answer is too given by jimthompson there.. :P

OpenStudy (anonymous):

So, now we are both equal.. Be happy.. :P

OpenStudy (mayankdevnani):

@ganeshie8

OpenStudy (mayankdevnani):

lolz.. @waterineyes

OpenStudy (mayankdevnani):

@ikram002p

OpenStudy (anonymous):

Yeah, surely we need experts' advice here.. :) I have got rusty on set related things.. :)

OpenStudy (mayankdevnani):

:P

OpenStudy (anonymous):

It seems like you have read it now somewhere and then you have come here.. :P

OpenStudy (mayankdevnani):

i get it !!! Think @waterineyes \[\large \bf [-1,0,1]\]

OpenStudy (anonymous):

Okay, I am thinking.. @mayankdevnani

OpenStudy (mayankdevnani):

if we choose 1 from the set,now there are two numbers remaining i.e -1 and 0. That's why,we can't add 1+1 ???

OpenStudy (anonymous):

yes?? I am in doubt.. :)

OpenStudy (anonymous):

you have to select two numbers from the set, why can't I choose one number twice?

OpenStudy (ikram002p):

for part A 1+1=2 for B 1/0 = :O for D -1-1=-2 so if i were u i would only chose C O.O

OpenStudy (anonymous):

yippie, Ikram is on my side.. :P

OpenStudy (mayankdevnani):

lolzz

OpenStudy (mayankdevnani):

but JIMTHOMPSON

OpenStudy (mayankdevnani):

he is on my SIDE

OpenStudy (mayankdevnani):

:)

OpenStudy (anonymous):

Sorry talk about online members.. :P

OpenStudy (ikram002p):

hmmm im on my side too ^_^

OpenStudy (mayankdevnani):

GREAT ^^^

OpenStudy (anonymous):

That does not matter here.. :P

ganeshie8 (ganeshie8):

i would go with A, B, C lets conduct a poll maybe :)

OpenStudy (anonymous):

B also??

OpenStudy (mayankdevnani):

`poll` Whow !!!!

OpenStudy (anonymous):

Sorry, but you are alone here @ganeshie8 :P

OpenStudy (mayankdevnani):

@ganeshie8 you can't go with option B because 1/0=undefined which is open in the interval

OpenStudy (anonymous):

One thing I am sure about, ganesh has proven that also as wrong.. I was sure about Division.. What man.!! :P

ganeshie8 (ganeshie8):

thats precicely the reason we can go with it, 1/0 is not defined in any of the number sets yet the real numbers are closed under division and multiplication right ?

ganeshie8 (ganeshie8):

since 1/0 is not defined, we don't bother about it when considering closure

OpenStudy (ikram002p):

ganesh R* is closed :) R itself nope

OpenStudy (anonymous):

No ganesh.. :)

ganeshie8 (ganeshie8):

then il need updates :) whats R* ikram ?

OpenStudy (ikram002p):

R*=R-{0}

OpenStudy (mayankdevnani):

@ganeshie8 think that :- the set is closed it means that only -1,0,1 (integral values) are considered. 1/0=undefined and UNDEFINED is not in this particular set !!!

OpenStudy (mayankdevnani):

please think over this

OpenStudy (mayankdevnani):

*UNDEFINED is not considered in this set

OpenStudy (ikram002p):

only under division R is not closed hmm but this is also an axiom thingy :P some books dont mention that so its ok sometimes to say R is closed on division

OpenStudy (mayankdevnani):

no,we can't say that UNDEFINED is closed because its UNDERSTOOD or apply common sense,how can we arrange UNDEFINED ??

OpenStudy (mayankdevnani):

@ikram002p

ganeshie8 (ganeshie8):

okay tjat makes sense, so real numbers are not closed under division because 1/0 is undefined ?

OpenStudy (mayankdevnani):

If i would say that :- arrange planets in the market for market sale Now can you arrange real planets for MARKET SALE

ganeshie8 (ganeshie8):

this guy failed me again http://mathforum.org/library/drmath/view/61270.html

OpenStudy (ikram002p):

hehe yeah was gonna say about never heard of subtraction or division xD but i thought explaining it like this more better :P

Miracrown (miracrown):

Ganeshie - Don't feel discouraged, that ''guy'' is a doctor for a reason. Hehe.

ganeshie8 (ganeshie8):

yeah i am still with him :) just trying to manage the strong opposition here haha!

OpenStudy (mayankdevnani):

i think :- \[\large \bf ({- \infty,+ \infty})\] i think @real numbers can be arranged in this set,

OpenStudy (mayankdevnani):

don't you think about that ? @ganeshie8 @ikram002p

OpenStudy (ikram002p):

well i like to be simple xD i dont wanna show up hehe so IMHO its only C , otherwise hmm i dont wanna comment :P

ganeshie8 (ganeshie8):

integers are closed under division makes sense because 1/3 is defined yet not an integer however 1/0 itself is undefined, so real numbers are closed under division makes more sense to me..

ganeshie8 (ganeshie8):

but i believe u more @ikram002p so lets kick option B

OpenStudy (mayankdevnani):

yeah,,,option C would definitely satisfies. option C is one of the answers

OpenStudy (ikram002p):

ganesh integer is not closed under division or multiplication this is what u trying to say right ?

ganeshie8 (ganeshie8):

integers are closed under multiplication right ?

OpenStudy (ikram002p):

nope

ganeshie8 (ganeshie8):

ugh, let me think

OpenStudy (ikram002p):

there is no such a^-1 in N for all a in N such that aa^-1=1 :O

OpenStudy (mayankdevnani):

@ikram002p and @ganeshie8 scroll up and see those comments from which the doubt has occur ? and please be patient while you read all the comments and if you read all that comments,then please reply !!!

OpenStudy (mayankdevnani):

please TAKE YOUR TIME !!!

OpenStudy (mayankdevnani):

*while you are reading

ganeshie8 (ganeshie8):

how is multiplcation even related to inverse ?